OpenStudy (helenalunch):

-6 = 15 + x (CHECK MY ANSWER!) Please :)

OpenStudy (helenalunch):

-6 = 15 + x

OpenStudy (helenalunch):

so using inverse im subtracting 15 from 15 and -6

OpenStudy (helenalunch):

15-15 gives me x and then -6 minus 15 im getting 9 therefore 9 = x

OpenStudy (sunnnystrong):

Hmm.... 

OpenStudy (helenalunch):

This confuses me because the "study problem" we received was -9 = 19 + x told to subtract 19 from both 19 and -9 19-19 = x - 9 minus 19 gave 10 therefore x = 10 solved by the teacher themselves

OpenStudy (wanell):

It's -21

OpenStudy (sunnnystrong):

@helenalunch ... -9-19 does not equal 10 it's -28. Another way to check and see if your arithmetic is correct is by plugging x into the original equation and seeing if it's correct

OpenStudy (helenalunch):

That is literally the problem and answer inside the textbook and the explanation given by the teacher though. This is multiple transformations aswell as using inverse operations.

OpenStudy (sunnnystrong):

@wanell ... it's 21.

OpenStudy (sunnnystrong):

@helenalunch ... welll it doesn't change the fact that -19-9=-28 & not 10. Checking your first answer & your work... your subtracting incorrectly.

OpenStudy (wanell):

-21 check your answer

OpenStudy (sunnnystrong):

@wanell ooops copied the problem down wrong but same process. you should show your work versus giving direct answers.

OpenStudy (wanell):

well i could but its a little to late for that

OpenStudy (sunnnystrong):

\[-6=15+x\] *Subtract 15 from both sides \[-21=x\]